OpenStudy (anonymous):

y'' +4y' +4y = e^(3t) using variation of parameters I think I did it wrong because I got W(y1,y2) = 0

OpenStudy (anonymous):

Variation of parameters says if you have a homogenous solution: \[y_H(t)=y_1(t)+y_2(t)\] Then the solution is: \[y(t)=-y_1 \int\limits y_2 g(t)/W(t)dt +y_2 \int\limits y_1 g(t)/W(t)dt + y_H(t)\] So: \[y_H(t) = c_1 e^{-2t} + c_2 t e^{-2t}\] Because of the redundancy in the roots right?

OpenStudy (anonymous):

I get W(t)=e^(-4t)

OpenStudy (anonymous):

Does all that seem good?

OpenStudy (anonymous):

W(t)= y1 x y'2 - y2 x y'1

OpenStudy (anonymous):

i used this and i get: y1 x y'2 - y2 x y'1 e^(-2t) x -2e^(-2t) -e^(-2t) x -2e^(-2t) = 0? that is where i get 0

OpenStudy (mendicant_bias):

I get the particular solution cutting itself down to:\[\frac {e^{3x}}{25}\] Anybody else got this?